OpenStudy (anonymous):

pipe A is a inlet pipe filling the tank at 8000l/hr and pipe B is the outlet pipe which empties the tank in 3 hours.the capacity of the tank is

OpenStudy (anonymous):

pipe A=1/x, pipe B=1/3, so 8000=1/x-1/3

OpenStudy (anonymous):

@waterineyes

OpenStudy (perl):

does that say 8000 liters / hr ?

OpenStudy (anonymous):

yes

OpenStudy (perl):

so you want to know how big the tank is , in liters

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Congratulations @perl for going \(\color{green}{\textsf{Green..}}\)

OpenStudy (perl):

can you double check your question for typos. something is a bit odd

OpenStudy (perl):

does it give you the rate of the outlet pipe?

OpenStudy (anonymous):

nope

OpenStudy (perl):

can you take a picture and upload the question

OpenStudy (anonymous):

yes this bit tough question

OpenStudy (perl):

i would like to see your worksheet

OpenStudy (anonymous):

pipe A fill at 8000liters/hour pipe B empties at xliters/hour

OpenStudy (anonymous):

1/x=pipe A,1/3=pipe B

OpenStudy (perl):

be careful with your units. if you multiply x liters / hr * 3 hr = total capacity in liters

OpenStudy (anonymous):

8000*t=x*3

OpenStudy (perl):

that has two unknowns , you cant solve it

OpenStudy (anonymous):

8000t=3x, so x=8000t/3=x

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

@ganeshie8

OpenStudy (phi):

I would say fill rate * time = Volume If you can empty the tank in 3 hours using pipe B, the rate (out) is rate * 3 = V rate= V/3 if you fill the tank at a rate of 8000 L/hr using pipe A you say T * 8000 = V we have two unknowns : V and T we need more info

OpenStudy (anonymous):

please anybody try to solve or explain .don't say two unknowns

OpenStudy (anonymous):

yes i agree @phi.

OpenStudy (phi):

the solution would be a line V= 8000 T for example, if T= 1 hour, then V is 8000 L pipe B empties this in 3 hours , so its rate would be 8000/3 or, if T=2 V = 16000 L and B's rate is 16000/3

OpenStudy (anonymous):

i have options for the question a)12000 b)8000 c)6000 d)4000

OpenStudy (phi):

Is there more info ? For example, does pipe B take 3 hours to empty the tank at the same time that 8000 L is flowing in from pipe A ?